# Forecasting Response to Treatment with Global Deep Learning and Patient-Specific Pharmacokinetic Priors

### v6

@@ B.3 Model MAE and RMSE results
Baseline Exogenous Baseline Exogenous
ETS 10.040 13.531
@@
Table 6: Root Mean Square error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data.
@@
Baseline Exogenous Baseline Exogenous
ETS 15.754 23.151
@@
Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data. Average results are computed across 8 trials.
Model All Values Critical Values
Baseline Exogenous Baseline Exogenous
ETS 10.126 11.781
@@
MLP 10.568 10.275 11.805 11.434
(0.628) (0.183) (0.731) (0.186)
LSTM 10.604 9.606 12.999 11.144
(0.929) (0.372) (2.294) (0.708)
TCN 11.011 11.090 12.693 13.212
(1.066) (0.708) (1.334) (0.886)
TFT 9.605 9.515 10.686 10.555
(0.735) (0.310) (0.986) (0.495)
NHITS 9.234 8.954 10.436 10.117
(0.056) (0.058) (0.125) (0.086)
NHITS-Sum Total 8.988 10.092
(0.023) (0.082)
NHITS-Pharmacokinetic 8.769 10.007
(0.069) (0.087)
TFT-Pharmacokinetic 9.492 10.484
(0.399) (0.374)
Table 8: Root mean square error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data.
Model All Values Critical Values
Baseline Exogenous Baseline Exogenous
ETS 15.351 17.478
(0.00) (0.00)
MLP 16.128 15.625 17.942 17.482
(0.580) (0.131) (0.716) (0.232)
LSTM 15.766 14.704 18.565 16.784
(0.930) (0.297) (1.818) (0.688)
TCN 16.173 16.192 18.432 18.978
(0.932) (0.545) (1.097) (0.806)
TFT 15.031 14.820 16.720 16.455
(0.576) (0.186) (0.954) (0.413)
NHITS 15.106 14.596 16.901 16.421
(0.075) (0.080) (0.168) (0.133)
NHITS-Sum Total 14.714 16.438
(0.043) (0.135)
NHITS-Pharmacokinetic 14.375 16.287
(0.126) (0.154)
TFT-Pharmacokinetic 14.846 16.351
(0.366) (0.284)

### v7

@@ C.5 Model MAE and RMSE Results
Baseline Exog. Baseline Exog.
MAE
ETS 9.450 13.457
@@
MLP 15.561 11.461 22.380 14.449
(2.318) (0.553) (3.568) (0.502)
RNN 19.033 27.116 28.834 43.256
(3.479) (12.484) (6.974) (23.190)
LSTM 18.567 13.723 26.863 18.028
(11.191) (2.373) (18.642) (3.661)
TCN 17.094 13.551 25.163 16.908
(1.719) (0.628) (3.092) (1.074)
NHITS 12.918 10.546 18.436 13.371
(0.064) (0.106) (0.369) (0.387)
NHITS-ST 10.082 12.679
(0.054) (0.135)
NHITS-PK 10.039 12.339
(0.188) (0.435)
Table 7: Mean absolute error (MAE) (top) and Root Mean Square Error (RMSE) (bottom) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data. Average results are computed across 8 trials.
@@
Baseline Exog. Baseline Exog.
MAE
ETS 9.022 10.028
@@
NHITS-PK 14.385 16.251